A book club offers a choice of 3 books from a list of 30. How many different selections of 3 different books each can be made from this list?
I TRIED 30*29*28= 24360 THEN DIVIDED THAT BY 3 AND GOT 8120
BUT IT WAS STILL WRONG.

Answer by Alan3354(69443)

A book club offers a choice of 3 books from a list of 30. How many different selections of 3 different books each can be made from this list?
-------------
The 1st choice is 1 of 30, then 29, then 28
--> 30*29*28 = 24360
----------
Choosing A, B & C is the same as A, C & B or C, B & A, etc.
There are 6 possible arrangements for 3 books, 3*2*1 = 6
24360/6 = 4060 possibilities